OpenStudy (anonymous):

Solve: log(3) ^(x+4) - log(3) ^x =2

hartnn (hartnn):

use the log property \(\Large \log A -\log B = \log (A/B)\)

hartnn (hartnn):

is it \(\Large \log 3^{x+4}\) or \(\Large \log_3 (x+4)\)

OpenStudy (anonymous):

\[ \log(3) ^{(x+4)} - \log(3) ^x =2\]\[(x+4) \log(3) - (x)\log(3) =2\]\[4\log(3) =2\] So I would therefore sy it is the second thing hartnn posted.

OpenStudy (anonymous):

\[\log_3x+4-\log_3x=2\] like tus right ?

hartnn (hartnn):

yup, that makes more sense

OpenStudy (anonymous):

Yes, it is log(3) x+4, sorry it's to differentiate on my paper

hartnn (hartnn):

ok, use this property for left side \(\Large \log A -\log B = \log (A/B)\) what do u get ?

OpenStudy (anonymous):

log(3) (x+4)/x

hartnn (hartnn):

correct!

hartnn (hartnn):

\(\log_3 [(x+4)/x] = 2\) now use the definition \(\Large \log _AB=C \implies B=A^C\)

OpenStudy (anonymous):

I'm not sure I'm following that one

OpenStudy (anonymous):

Like, 3^2= (x+4)/x?

hartnn (hartnn):

yeah, thats correct! :) good.

OpenStudy (anonymous):

Then solve for x?

hartnn (hartnn):

now can you solve further ? 3^2 = 9 then multiply both sides by x

OpenStudy (anonymous):

9x=x+4 8x=4 x=2 It's saying the correct answer is 1/2

OpenStudy (anonymous):

\[\log_3x+4-\log_3x= 2\]\[\log_3x+4-\log_3x=2~ \log_33\]\[\log_3x+4-\log_3x=\log_33^{2}\]\[\log_3x+4-\log_3x=\log_39\]\[\log_3x+4=\log_39+\log_3x\]\[\log_3x+4=\log_39x\]\[x+4=9x\]\[4=8x\]\[x=1/2\]

OpenStudy (anonymous):

8x=4 not 4x=8

hartnn (hartnn):

8x = 4 gives x= 4/8 = 1/2

OpenStudy (anonymous):

Oh, I see. Thank you!

hartnn (hartnn):

thats a nice alternative @WHAT?! , good!

OpenStudy (anonymous):

tnx :) I just don't like fractions inside the logs.